The K base β X-ray of argon has a wavelength of 0.36 nm. The minimum energy needed to ionize an argon atom is 16 eV. Find the energy needed to knock out an electron from the K shell of an argon atom.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
λ = 0.36 nm E = 1242/0.36 = 2450 eV (E base M – E base K) Energy needed to ionize an organ atom = 16 eV Energy needed to knock out an electron from K-shell = (3450 + 16) eV = 3466 eV = 3.466 KeV.